OpenStudy (anonymous):

what is the solution to this system of linear equations? y=-6x+52 y=2x+17

OpenStudy (anonymous):

Well there are multiple ways of solving this system of equations. The one I'm going to use in this case will be to just set them equal to each other. y=-6x+52 y=2x+17 (tip is to think that y is the same thing as y. Like an apple is the same thing as an apple. So if y equals both of those, then when you finally get x, they'll end up being the same thing.) 2x+17=-6x+52 (then just solve for x) +6x +6x 8x+17=52 -17 -17 8x=35 /8 /8 x=4.375 From here, to find y, just plug the solution for x into one of your equations. I'm just going to do y=2x+17 because it's a little easier than the other. y=2x+17 y=2(4.375)+17 y=8.75+17 y=25.75 Not the prettiest number, but let's check now by plugging in both the solutions for x and y into one of the equations just to make sure. This time I'll just go with the first equation... just because I did the second one last time. I'm trying to spice it up here. y=-6x+52 25.75=-6(4.375)+52 25.75=-26.25+52 25.75=25.75 Yeah! We did it!! So your answer is... y=25.75 x=4.375 but your teacher will probably want you to write it as... (4.375, 25.75)